# Latent Representation Learning of Multi-scale Thermophysics: Application to Dynamics in Shocked Porous Energetic Material

### v1

@@ 1 Introduction
Multi-scale dynamics, coupling phenomena at two or more distinct scales, arise in many physical applications, such as multi-phase flows (saurel1999multiphase, ; udaykumar1997multiphase, ), transport in porous media (mcdowell1986particle, ; parker1989multiphase, ), multi-material flows (luo2004computation, ; banks2007high, ), etc. When scales are distinctly separated, distinguished by the presence of an underlying microstructure, for example (perry2018relating, ), the conventional practice is to solve the dynamics at the observable (macro-) scale by homogenizing the properties and processes of the unresolved (micro-/meso-) scale. This leads to the need for closure models to connect physics across disparate scales.
Scale bridging has been accomplished in a variety of ways (fish2010multiscale, ; liu2006bridging, ; saurel2018multiscale, ); in recent times, data-driven closure models have been developed using highly resolved direct numerical simulations (DNS) to inform machine learning algorithms (sanderse2024scientific, ; beck2019deep, ; san2018neural, ). The machine learned models serve as surrogates to close macro-scale dynamics. With the exponential rise of deep learning (DL) (lecun2015deep, ; bengio2017deep, ), artificial intelligence (AI)/machine learning (ML) has received a great deal of research attention in the context of scale-bridging for multi-scale dynamics problems. In particular, physics-informed machine Learning (PIML) (karniadakis2021physics, ) approaches seek to assimilate sub-grid dynamics in deep neural networks (DNNs), and a wide range of capabilities for multi-scale modeling using these techniques are now available (nguyen2023parc, ; kashinath2021physics, ; oommen2024rethinking, ).
While these DNN-based approaches are powerful for data assimilation and representation, they are inherently data hungry, which places the onus on computationally intensive physical simulations to supply adequately rich datasets to the AI/ML model. In this paper, we propose a meta-learning framework in which one leverages an ensemble of inexpensive micro-scale simulations in order to accelerate the learning process of a surrogate model for the meso-scale dynamics.
Here, the meta-learning approach is applied to an archetypal multi-scale dynamics problem, namely the response of a porous energetic material (EM) to shock loading. EM (defining the broad class of propellants, explosives, and pyrotechnics) have complex microstructural morphologies, which strongly influence the performance characteristics of these materials (baer2002modeling, ). Defects such as voids and cracks in the microstructure of EM are sites of energy localization, which can evolve into concentrated regions of high temperature called “hotspots” (field1992hot, ). Under a strong shock condition, a major mechanism for the formation of hotspots is void collapse, which leads to energy release on a very short time-scale (i.e., nanoseconds) (menikoff2004pore, ; tarver1996critical, ).
@@
The problem of shock-induced energy localization in porous energetic material is inherently multi-scale in the following sense: at the micro-scale (i.e., $\mathcal{O}(1\,\mu m)$), the collapse of isolated voids in the microstructure of an EM can create localized hotspots that may expand and lead to ignition of the surrounding material; the resulting hotspots can trigger chemical reactions that release further energy and intensify the traveling shockwave, which, eventually, can lead to a self-sustaining detonation at the macro-scale (i.e., $\mathcal{O}(1\,mm)$). [^1] In the context of porous energetic material, we refer to the length scale of a single pore (i.e., $\mathcal{O}(1\,\mu m)$) and the length scale of a field of voids in a microstructure domain of size $\mathcal{O}(100\,\mu m)$ and $\mathcal{O}(1\,mm)$ as the micro, meso and macro-scale, respectively. For this to occur, however, the heat release from a large ensemble of hotspots must be coordinated with the overpassing shock, successively pumping energy into the shock all the way to the formation of a detonation wave, a phenomenon called shock-to-detonation transition (SDT) (handley2018understanding, ). Thus, the presence of multitudes of potential localization sites in the stochastic microstructure of porous energetic materials needs to be simulated to model SDT (bernecker1986deflagration, ).
@@
Direct numerical simulations (DNS) play an important role in understanding the above-mentioned complex physics at the meso-scale (kapahi2013dynamics, ; rai2015mesoscale, ; sen2018multi, ; seshadri2022meso, ; Nguyen2022multiscale, ). Figure 1 (a) and (b) illustrate an instance of numerical simulation results representing the shock-induced temporal evolution of temperature and pressure fields associated with a micro-scale single-pore collapse (nguyen2022multi, ) and a meso-scale field of voids in a domain of EM microstructure (seshadri2022meso, ), respectively. The columns in the figures show a snapshot of the corresponding temperature and pressure fields at different instants of time (noted above the panels); In addition, there has been a recent interest in data-driven approaches, especially deep learning models, to learn surrogate models from simulation data, and make fast predictions of the corresponding dynamics with comparable accuracy to DNS (nguyen2023parc, ). However, the high computational cost of DNS for generating meso-scale numerical simulations poses practical challenges in data-driven approaches, which typically need relatively large datasets for training the models. Alleviating the need for computationally intensive meso-scale simulations motivates the current meta-learning strategy.
@@
In general, the cross-disciplinary areas of meta-learning (hospedales2021meta, ; lee2022meta, ) and transfer-learning (pan2009survey, ; olivas2009handbook, ) consist of a class of machine learning methodologies that aim to learn a common “meta” information from a family of closely related tasks and use it to quickly adapt the learned model to new unseen tasks. We use an analogy between natural language processing (NLP) and the relation between micro- and meso-scale dynamics, associated with the response of microstructured material to external loads, in the following way: an important step in the learning process of an NLP model is the tokenization in which a piece of text, e.g., a sentence or a paragraph, is broken down into meaningful chunks of information, e.g., words, which are called tokens and can be considered as discrete building blocks of the original text. The tokenization lets an NLP model start by analyzing these building blocks, and in later steps, combine the token-level learned information to capture the correlation between tokens and learn higher-level information about the whole text.
@@
Following the above-mentioned tokenization analogy, we consider the meso-scale dynamics (“paragraph”) as a correlated system of machine learning building blocks, where each constituent encodes a reduced (latent) representation of the micro-scale dynamics (“words”). The main hypothesis underlying our consideration of a reduced representation of micro-scale dynamics, as building blocks for meso-scale, is that the fine-grain and stochastic nature of the micro-scale physics has a negligible effect on the coarse-grain meso-scale behavior of the system, and it mostly gets smeared out from the statistical point of view. Therefore, we consider the direct sum of the reduced representations of the micro-scale dynamics, after applying the appropriate time-delay, as a “first order approximation” of the meso-scale dynamics for a short time span. However, we assume that the interaction between these micro-scale building blocks would play a fundamental role in determining the meso-scale evolution of the corresponding physical fields. In other words, for longer time periods than the micro-scale temporal resolution, the evolution of physical fields over neighboring micro-scale building blocks will couple together, leading to deviation of the meso-scale dynamics from purely being the direct sum of micro-scale dynamics. The effects due to the correlation between neighboring micro-scale building blocks are considered as “higher order corrections.”
@@
Therefore, the proposed meta-learning framework can be summarized, roughly, into two steps: in the first step, one learns a reduced representation of the micro-scale dynamics, as the building blocks of a “first order approximation” of the meso-scale dynamics, from a relatively large dataset of (inexpensive) micro-scale numerical simulations; in the second step, one learns the correlation between these building blocks, “higher order corrections,” from a relatively small dataset of (expensive) meso-scale simulations. In this paper, we test the hypothesis underlying the proposed meta-learning framework for the problem of shock-induced energy localization in porous energetic materials.
@@
The structure of the paper is as follows. In Section 2, we provide an overview of the main ideas underlying our proposed approach, and explain how it can be connected to reduced order modeling and meta-learning. In Section 3, we present the detailed mathematical formulation and architecture of the model considered in this work. In Section 4, we present our proposed meta-learning framework applied to the problem of learning a surrogate model for the meso-scale dynamics associated with shock-induced energy localization in EM. Furthermore, we compare the prediction performance of our model with a physics-aware recurrent convolutional neural network (PARC), in the scarce meso-scale data regime, in terms of sensitivity quantities of interest (QoI) for EM. Also, the uncertainty in our model’s predicted QoI, due the stochastic nature of learned dynamics over the latent space, is investigated. We conclude in Section 5 with a summary of our work and potential future directions for expanding the current work. Additionally, we provide details on implementation of our model and its training in Appendix A.1. In Appendix A.2, we discuss more details on the numerical simulation setup considered for generating the data used in this work.
2 Overview and Related Works
The proposed approach in this work combines ideas from reduced order modeling (ROM) (quarteroni2014reduced, ) and meta-learning (hospedales2021meta, ) in machine learning to accelerate the learning process of surrogate models for the meso-scale dynamics underlying the response of microstructured material to external loads, based on the relation between micro and meso-scale dynamics. In the following, we review the main ideas and the contribution of this work.
Reduced Order Modeling
Almost all classical projection-based reduced order modeling methods, including Proper Orthogonal Decomposition (POD) (berkooz1993proper, ) and Dynamic Mode Decomposition (DMD) (schmid2010dynamic, ), are based on considering a projection onto a linear subspace of the full state space. However, the projection-based ROMs can only provide an accurate model in the case of problems with a fast-decaying Kolmogorov $n$-width, e.g., diffusion-dominated problems, whose solution space can be well approximated by a linear subspace (lee2020model, ). On the other hand, many physical problems of interest, including advection-dominated problems like shock-induced energy localization in EM, exhibit a slow-decaying Kolmogorov $n$-width, which necessitates the consideration of nonlinear techniques for approximating the solution space. Each of various nonlinear techniques in reduced order modeling try to approximate the invariant (slow) manifold (gorban2005invariant, ; carr2012applications, ) of the corresponding dynamical system which can be defined as an attracting submanifold of the system’s full state space on which the dynamics is slow and free of high-frequency motions compared to the state space regions away from its basins of attraction (see, e.g., (benner2015survey, ; fresca2021comprehensive, ; geelen2023operator, )). In particular, the deep learning-based reduced order models (DL-ROMs) have become popular recently for learning nonlinear latent representations of physical dynamics in deep neural networks (maulik2021reduced, ; kim2022fast, ; fries2022lasdi, ; vlachas2022multiscale, ; conti2023reduced, ; bonneville2024gplasdi, ). However, to the best of our knowledge, most of DL-ROMs aim to learn a deterministic dynamics over the invariant manifold of the system which limits the generalizability and robustness of the model.
In contrast, we consider a probabilistic deep learning-based model which learns a stochastic dynamics over the latent space. This stochasticity play a crucial role in our model’s success in learning the meso-scale dynamics of shock-induced energy localization in EM in scarce data regimes. Our model consists of two main parts. In the first part, we use variational autoencoders (VAE) (kingma2014auto, ) to encode the corresponding system of coupled physical fields, whose evolution describes the dynamics in the full state space, into discrete Gaussian random fields as the associated latent representations. In the second part of our model, we learn the reduced dynamics of the corresponding Gaussian random fields using an autoregressive model. Motivated by the stochastic differential equation (SDE) formulation, the autoregressive functional of our model is decomposed into two parts: a deterministic part that captures the evolution of the latent mean fields over the slow manifold; and a stochastic part that encapsulates the evolution of the latent variance fields as the probabilistic deviation from the slow deterministic dynamics. We further decompose each of the deterministic and the stochastic autoregressive models into two functionals: a decoupled part which captures the short time-span evolution of each latent field corresponding to a single physical field, independent of the other physical fields; and a correlation part which encapsulates the interaction between all the involved physical fields in our multiphysics system. This strategy allows for learning the long-term global dynamics of coupled physical fields, which can involve distinct phases over different time periods, such as the hotspot ignition and growth phases in shock-induced energy localization in EM.
Meta-Learning
Broadly speaking, meta-learning (schmidhuber1987evolutionary, ; bengio1990learning, ) or learning to learn (thrun1998learning, ) refers to a bi-level learning process in which an inner (base) learning algorithm solves a particular task defined by a dataset and objective function, while an outer (meta) algorithm updates the inner learning algorithm such that the learned model improves an outer objective like generalization over a family of closely related tasks. The meta-learning approaches can be roughly divided into three main classes (penwarden2023metalearning, ): i) methods based on a bi-level optimization, where the inner algorithm learns the parameters of the model conditioned on certain hyperparameters and these hyperparameters are optimized in the outer level based on a meta-loss (andrychowicz2016learning, ; ravi2017optimization, ; franceschi2018bilevel, ); ii) methods that learn an outer-level metric space which is based on a notion of distance (similarity) between inner-level tasks (koch2015siamese, ; vinyals2016matching, ; snell2017prototypical, ); iii) the feed-forward methods that take the hyperparameters and task datasets as input and predict the optimal model parameters (hochreiter2001learning, ; santoro2016meta, ; mishra2017simple, ). The application of meta-learning approaches to physics problems has been mostly in the context of physics-informed neural networks (PINNs) (raissi2019physics, ) to improve their performance over a parametrized family of partial differential equations (liu2022novel, ; psaros2022meta, ; bihlo2024improving, ) .
However, in this work, we leverage the relation between micro- and meso-scale dynamics in the response of microstructured material to external loads to propose a meta-learning framework for learning closure models at the meso-scale. Motivated by the idea of tokenization in natural language processing, we consider a meso-scale domain of microstructured material as a correlated system of interacting micro-scale building blocks. While an inner level model learns the micro-scale dynamics of each constituent independently, an outer level model “stitches” them together by learning the correlation between these micro-scale building blocks.
@@ 3 Method
Consider temporal partial differential equations (PDEs) of the form
$\begin{split}&\partial_{t}\mathbf{u}=\mathcal{F}\left({x,\mathbf{u},\partial_{x}\mathbf{u},\partial_{xx}\mathbf{u},\cdots;\mathfrak{a}}\right)\,,\qquad t\in[0,\mathbb{T}]\,,\;\;x\in\mathbb{X}\,,\\ &\mathbf{u}(0,x)=\mathbf{u}^{0}(x)\,,\qquad x\in\mathbb{X}\,,\\ &B[\mathbf{u}](t,x)=0\,,\qquad t\in[0,\mathbb{T}]\,,\;\;x\in\partial\mathbb{X}\,,\end{split}$ (3.1)
parametrized by a family of static parameters $\mathfrak{a}$. The nonlinear functional $\mathcal{F}$ in (3.1) encapsulates the governing dynamics of a system of coupled physical fields $\mathbf{u}=(u_{1},u_{2},\cdots,u_{k}):[0,\mathbb{T}]\times\mathbb{X}\to\mathbb{R}^{k}$ over the spatial domain $\mathbb{X}$ in time interval $[0,\mathbb{T}]$. The notations $\partial_{x}\mathbf{u}$ and $\partial_{xx}\mathbf{u}$ denote the first- $\frac{\partial\mathbf{u}}{\partial x}$ and second-order $\frac{\partial^{2}\mathbf{u}}{\partial x^{2}}$ partial derivatives w.r.t. the spatial dimension. In addition, $\mathbf{u}^{0}(x)$ specifies the initial condition and $B[\mathbf{u}](t,x)=0$ gives the boundary conditions over the boundary $\partial\mathbb{X}$ of the spatial domain $\mathbb{X}$. The parameters $\psi=(\mathfrak{a},\partial\mathbb{X})\in\Psi$ represent physical and geometrical properties of the system, e.g., the material properties and the global shape of the boundary of the spatial domain.
@@
of the state vector $\mathbf{U}$ corresponding to various initial conditions and parameter values, we want to learn a surrogate model, which can emulate the time evolution of the dynamical system (3.2). We assume that the dominant part of the dynamics of the system happens over a lower-dimensional invariant manifold (slow manifold) embedded inside the high-dimensional input space $\mathbb{R}^{k\times N}$ (temam2012infinite, ; holmes2012turbulence, ). Hence, we focus on learning a rich enough latent representation of the dynamical system and the corresponding evolution over the latent space.

### v2

@@ 1 Introduction
Multi-scale dynamics, coupling phenomena at two or more distinct scales, arise in many physical applications, such as multi-phase flows (saurel1999multiphase, ; udaykumar1997multiphase, ), transport in porous media (mcdowell1986particle, ; parker1989multiphase, ), multi-material flows (luo2004computation, ; banks2007high, ), etc. When scales are distinctly separated, distinguished by the presence of an underlying microstructure, for example (perry2018relating, ), the conventional practice is to solve the dynamics at the observable (macro-) scale by homogenizing the properties and processes at the unresolved (micro-/meso-) scale. This leads to the need for closure models to connect physics across disparate scales.
@@
The problem of shock-induced energy localization in porous energetic material is inherently multi-scale in the following sense: at the micro-scale (i.e., $\mathcal{O}(1\,\mu m)$), the collapse of isolated voids in the microstructure of an EM can create localized hotspots that may expand and lead to ignition of the surrounding material; the resulting hotspots can trigger chemical reactions that release further energy and intensify the traveling shockwave, which, eventually, can lead to a self-sustaining detonation at the macro-scale (i.e., $\mathcal{O}(1\,cm)$). [^1] In the context of porous energetic material, we refer to the length scale of a single pore (i.e., $\mathcal{O}(1\,\mu m)$) and the length scale of a field of voids in a microstructure domain of size $\mathcal{O}(1\,mm)$ and $\mathcal{O}(1\,cm)$ as the micro, meso and macro-scale, respectively. For this to occur, however, the heat release from a large ensemble of hotspots must be coordinated with the overpassing shock, successively pumping energy into the shock all the way to the formation of a detonation wave, a phenomenon called shock-to-detonation transition (SDT) (handley2018understanding, ). Thus, the presence of multitudes of potential localization sites in the stochastic microstructure (i.e., the meso-scale at order of 1mm1\,$m$) of porous energetic materials needs to be simulated to model SDT (bernecker1986deflagration, ).
@@
Direct numerical simulations (DNS) play an important role in understanding the above-mentioned complex physics at the meso-scale, where defects (pores, cracks) and interfaces in the heterogenegous microstructure play key roles in energy localization (kapahi2013dynamics, ; rai2015mesoscale, ; sen2018multi, ; seshadri2022meso, ; Nguyen2022multiscale, ). Figure 1 (a) and (b) illustrate an instance of numerical simulation results representing the shock-induced temporal evolution of temperature and pressure fields associated with a micro-scale single-pore collapse (nguyen2022multi, ) and a meso-scale field of voids in a domain of EM microstructure (seshadri2022meso, ), respectively. The columns in the figures show a snapshot of the corresponding temperature and pressure fields at different instants of time (noted above the panels). However, with current computational resources, fully resolved DNS simulations are not feasible for prediction and design in large, i.e., macro-scale, systems that employ energetic materials.
To enable design and predictive modeling, there has been increasing interest in multi-scale modeling with data-driven approaches, to make fast predictions of the macro-scale dynamics with comparable accuracy to DNS (nguyen2023parc, ). Scale bridging between micro- and macro-scales has been accomplished in a variety of ways (fish2010multiscale, ; liu2006bridging, ; saurel2018multiscale, ); in recent times, data-driven closure models have been developed using highly resolved direct numerical simulations (DNS) to inform machine learning algorithms (sanderse2024scientific, ; beck2019deep, ; san2018neural, ). The machine learned models serve as surrogates to close macro-scale dynamics. With the exponential rise of deep learning (DL) (lecun2015deep, ; bengio2017deep, ), artificial intelligence (AI)/machine learning (ML) has received a great deal of research attention in the context of scale-bridging for multi-scale, multiphysics problems. In particular, physics-informed machine Learning (PIML) (karniadakis2021physics, ) approaches seek to assimilate sub-grid dynamics in deep neural networks (DNNs), and a wide range of capabilities for multi-scale modeling using these techniques are now available (nguyen2023parc, ; kashinath2021physics, ; oommen2024rethinking, ). While DNN-based approaches are powerful for data assimilation and representation, they are inherently data hungry, which in turn places the onus on computationally intensive simulations to supply adequately rich datasets to the AI/ML model. In this regard, the high computational cost of DNS for generating meso-scale numerical simulations continues to pose practical challenges in data-driven approaches.
@@
Alleviating this need for computationally intensive meso-scale simulations motivates the current paper. Here, we propose a meta-learning framework in which one leverages an ensemble of inexpensive micro-scale (inexpensive single pore) simulations in order to accelerate the learning process of a surrogate model for the meso-scale (microstructure-level) dynamics. Broadly speaking, the cross-disciplinary areas of meta-learning (schmidhuber1987evolutionary, ; bengio1990learning, ; hospedales2021meta, ; lee2022meta, ) and transfer-learning (thrun1998learning, ; pan2009survey, ; olivas2009handbook, ) consist of a class of machine learning methodologies that aim to learn a common “meta” information from a family of closely related tasks and use it to quickly adapt the learned model to new unseen tasks. These methods are based on a bi-level learning process in which an inner (base) learning algorithm solves a particular task defined by a dataset and objective function, while an outer (meta) algorithm updates the inner learning algorithm such that the learned model improves an outer objective/physical solution. In this regard, we use an analogy in natural language processing (NLP) in the following way: an important step in the learning process of an NLP model is tokenization, in which a piece of text, e.g., a sentence or a paragraph, is broken down into meaningful chunks of information, e.g., words, which are called tokens and can be considered as discrete building blocks of the original text. Tokenization lets an NLP model start by analyzing these building blocks, and in later steps, combine the token-level learned information to capture the correlation between tokens and learn higher-level information about larger segments of text, such as sentences and paragraphs.
@@
Following this tokenization analogy, we consider the meso-scale dynamics (“paragraphs”) as a correlated system of machine learning building blocks, where each constituent encodes a reduced (latent) representation of the micro-scale dynamics (“words”). Therefore, we consider the direct sum of the reduced representations of the micro-scale dynamics (single pore collapse simulations as "words"), after applying an appropriate time-delay, as a “first order approximation” of the meso-scale dynamics for a short time span. However, for longer time periods and over larger spatial extents than the micro-scale, the physical fields over neighboring micro-scale building blocks will couple together (akin to interactions between words for constructing coherent “paragraphs”), so that the resulting meso-scale dynamics is not purely the direct sum of micro-scale dynamics. The effects due to interaction between neighboring micro-scale building blocks are considered as “higher order corrections.” In other words, we consider a decomposition of the meso-scale dynamics in the following (heuristic) form:
$\text{meso-dynamics}=\left(\bigoplus_{\text{micro-blocks}}\text{micro-dynamics}\times{\Delta t}_{\text{micro}}\right)+\left(\text{correlation terms}\times{\Delta t}_{\text{meso}}\right)\,$ (1.1)
@@
where ${\Delta t}_{\text{micro}}$ and ${\Delta t}_{\text{meso}}$ represent the micro- and meso-scale time-periods, respectively (see Section 3.2 for a more rigorous discussion). Therefore, the proposed meta-learning framework can be summarized into two steps: in the first step, one learns a reduced representation of the micro-scale dynamics, as the building blocks of a “first order approximation” of the meso-scale dynamics, from a relatively large dataset of (inexpensive) micro-scale numerical simulations; in the second step, by transferring the learned reduced representation of the micro-scale dynamics to a meso-scale domain, it suffices to learn the correlation between the temporal evolution of micro-scale building blocks. This second step needs training over a relatively small dataset of (expensive) meso-scale simulations compared to directly learning a surrogate meso-scale model from scratch.
The approach studied in this work combines ideas from the two research fields of meta-learning (hochreiter2001learning, ; koch2015siamese, ; andrychowicz2016learning, ; vinyals2016matching, ; santoro2016meta, ; ravi2017optimization, ; franceschi2018bilevel, ) and reduced order modeling (ROM) (berkooz1993proper, ; gorban2005invariant, ; schmid2010dynamic, ; carr2012applications, ; quarteroni2014reduced, ; lee2020model, ). To our knowledge, the application of meta-learning approaches to physics problems has been mostly in the context of physics-informed neural networks (PINNs) (raissi2019physics, ) to improve their performance over a parametrized family of partial differential equations (liu2022novel, ; psaros2022meta, ; bihlo2024improving, ). In addition, the proposed model architecture in this work, for learning a reduced representation of the micro-scale dynamics, is based on a probabilistic formulation of the latent dynamics in contrast to the most of deep learning-based ROMs which consider a deterministic latent dynamics (maulik2021reduced, ; kim2022fast, ; fries2022lasdi, ; vlachas2022multiscale, ; conti2023reduced, ; bonneville2024gplasdi, ). This probabilistic formulation leads to better generalizability and robustness of our model in scarce meso-scale data regime in addition to facilitating uncertainty quantification (UQ) in the prediction results.
@@
The structure of the paper is as follows. In Section 2, we present the detailed mathematical formulation and architecture of the model considered in this work. In Section 3, we present our proposed meta-learning framework applied to the problem of learning a surrogate model for the meso-scale dynamics associated with shock-induced energy localization in EM. Furthermore, we compare the prediction performance of our model with a physics-aware recurrent convolutional neural network (PARC), in the scarce meso-scale data regime, in terms of sensitivity quantities of interest (QoI) for EM. Also, the uncertainty in our model’s predicted QoI, due the stochastic nature of learned dynamics over the latent space, is investigated. We conclude in Section 4 with a summary of our work and potential future directions for expanding the current work. Additionally, we provide details on implementation of our model and its training in Appendix A.1. In Appendix A.2, we discuss more details on the numerical simulation setup considered for generating the data used in this work. Complementary results on the predicted evolution of the temperature $T$ and pressure $P$ fields by our model and PARC for additional test samples are presented in Appendix A.3.
@@ 2 Method
Consider partial differential equations (PDEs) of the form
[…]
parametrized by a family of static parameters $\mathfrak{a}$. The nonlinear functional $\mathcal{F}$ in (2.1) encapsulates the governing dynamics of a system of coupled physical fields $\mathbf{u}=(u_{1},u_{2},\cdots,u_{k}):[0,\mathbb{T}]\times\mathbb{X}\to\mathbb{R}^{k}$ over the spatial domain $\mathbb{X}$ in time interval $0,\mathbb{T}]$. The notations $\partial_{x}\mathbf{u}$ and $\partial_{xx}\mathbf{u}$ denote the first- $\frac{\partial\mathbf{u}}{\partial x}$ and second-order $\frac{\partial^{2}\mathbf{u}}{\partial x^{2}}$ partial derivatives w.r.t. the spatial dimension. In addition, $\mathbf{u}^{0}(x)$ specifies the initial condition and $B[\mathbf{u}](t,x)=0$ gives the boundary conditions over the boundary $\partial\mathbb{X}$ of the spatial domain $\mathbb{X}$. The parameters $\psi=(\mathfrak{a},\partial\mathbb{X})\in\Psi$ represent physical and geometrical properties of the system, e.g., the material properties and the global shape of the boundary of the spatial domain.
@@
of the state vector $\mathbf{U}$ corresponding to various initial conditions and parameter values, we want to learn a surrogate model, which can emulate the time evolution of the dynamical system (2.2). We assume that the dominant part of the dynamics of the system happens over a lower-dimensional invariant manifold (slow manifold) embedded inside the high-dimensional input space $\mathbb{R}^{k\times N}$ (temam2012infinite, ; holmes2012turbulence, ). Recall that slow manifold (gorban2005invariant, ; carr2012applications, ) of a dynamical system is defined as an attracting submanifold of the system’s full state space on which the dynamics is slow and free of high-frequency motions compared to the state space regions away from its basins of attraction (see, e.g., (benner2015survey, ; fresca2021comprehensive, ; geelen2023operator, )). Hence, we focus on learning a rich enough latent representation of the dynamical system and the corresponding evolution over the latent space.